# Revisiting Euclidean Alignment for Transfer Learning in EEG-Based Brain-Computer Interfaces

### v1

@@ Abstract
Due to the non-stationarity and large individual differences of EEG signals, EEG-based brain-computer interfaces (BCIs) usually need subject-specific calibration to tailor the decoding algorithm for each new subject, which is time-consuming and user-unfriendly, hindering their real-world applications. Transfer learning (TL) has been extensively used to expedite the calibration, by making use of EEG data from other subjects/sessions. An important consideration in TL for EEG-based BCIs is to reduce the data distribution discrepancies among different subjects/session, to avoid negative transfer. Euclidean alignment (EA) was proposed in 2020 to address this challenge. Numerous experiments from 10 different BCI paradigms demonstrated its effectiveness and efficiency. This paper revisits the EA, explaining its procedure and correct usage, introducing its applications and extensions, and pointing out potential new research directions. It should be very helpful to BCI researchers, especially those who are working on EEG signal decoding.
Keywords: Brain-computer interface, EEG, Euclidean alignment, label alignment, transfer learning
1 Introduction
@@
According to the input signals, BCIs can be broadly classified into invasive ones and non-invasive ones. The former use surgery to implant electrodes into the brain, which are mainly used for patients. Non-invasive BCIs do not need surgery, and hence are more convenient and preferred by able-bodied users. EEG is the most frequently used input signal for non-invasive BCIs.
@@
EEG signals are non-stationary, e.g., the same subject may have different EEG responses to the same stimulus in different sessions, and exhibit significant individual differences, i.e., different subjects have significantly different EEG responses to the same stimulus. As a result, it is very challenging, if not impossible, to design an EEG-based non-invasive BCI system that is plug-and-play and fits all subjects. Usually, for each new subject, we need to collect some subject-specific calibration data to tailor the parameters of the decoding algorithm, which is time-consuming and user-unfriendly.
@@
Transfer learning (TL), which utilizes data/knowledge from other subjects (sessions) to facilitate the calibration for a new subject (session), has been extensively used in EEG-based BCIs to expedite the calibration [11]. An important consideration in TL is how to reduce the data distribution discrepancies among different subjects (sessions) to avoid negative transfer [12].
@@
Euclidean Alignment (EA) [13] was proposed in 2020 to address this challenge. It uses two simple, efficient and closed-form formulas to align the raw EEG trials from different subjects/sessions, greatly facilitating TL [14]. It has been cited 261 times in WoS and 373 times in Google Scholar (as of 2/9/2025), ranking 2/344 of all papers published in IEEE Transactions on Biomedical Engineering [^2] https://webofscience.clarivate.cn/wos/alldb/summary/fd21b6a6-bdf1-4dcd-bae6-45d4041b195f-014738bfd9/times-cited-descending/1 (TMBE) in 2020, and 3/1810 of all papers published in IEEE TBME [^3] https://webofscience.clarivate.cn/wos/alldb/summary/5e219ca8-1532-4637-a27d-6c70f9cbbca2-014738a148/times-cited-descending/1 between 2020 and 2025. It was one of the only two Highly Cited Papers of IEEE TBME in 2020. It was also the number one contributor to the 2022 impact factor (4.6) of IEEE TBME [^4] https://jcr.clarivate.com/jcr-jp/journal-profile?journal=IEEE%20T%20BIO-MED%20ENG&year=2022&fromPage=%2Fjcr%2Fhome, among all 695 citable items published in 2020 and 2021.
EA has achieved remarkable performance in BCI competitions. By integrating it into our TL algorithms, the author’s team won nine National Championships in China BCI Competition (the most prestigious BCI competition in China, organized jointly by the National Natural Science Foundation of China, Chinese Institute of Electronics, and Tsinghua University) between 2019 and 2024. In the Benchmarks for EEG Transfer Learning [^5] https://beetl.ai/challenge (BEETL) competition organized at NeurIPS 2021, the champion team’s approach [15]for improving subject-independence is to align the latent feature distributions of the deep learning models between different subjects and sessions. This is similar to the Euclidean Alignment method, which performs spatial whitening with the average covariance matrix per subject, although the methods are not constrained to be applied on the model input.” In the runner-up’s approach [15], “Euclidean Alignment is used to close the gap between subjects in source domains and target domain. All trials of each subject are aligned such that the mean covariance matrices of all subjects are equal to the identity matrix after alignment. After alignment, data set distribution in both source domains and target domain are more similar.
@@
After five years of the publication of the EA paper [13], we now have deeper and more comprehensive understanding on how to properly use EA. Its applicable BCI paradigms have also been greatly expanded. This paper revisits the EA approach, explaining its procedure and correct usage, introducing its applications and extensions, and pointing out potential new research directions. It should be very helpful to BCI researchers, especially those who are working on EEG signal decoding.
@@
The remainder of this paper is organized as follows: Section 2 introduces the procedure, correct usage, and applications of EA. Section 3 introduces an extension of EA. Section 4 points out potential future research directions. Finally, Section 5 draws conclusions.
2 Euclidean Alignment (EA)
@@
This section introduces the EA approach [13]. It was inspired by Riemannian Alignment [16] proposed in 2018 in IEEE TBME, which has been cited 258 times in WoS and 370 times in Google Scholar (as of 2/9/2025) and is one of the five Highly Cited Papers of IEEE TBME in 2018.
2.1 The EA Algorithm
@@
The goal of EA is to make EEG data distributions from different subjects (sessions) more consistent (i.e., to reduce the inter-subject/session discrepancies), facilitating TL. In fact, after EA, often the EEG data from auxiliary subjects/sessions (called source domains in TL) can be combined directly with the limited amount of calibration data from the new subject/session (called target domain in TL) to train a classifier; the performance is usually quite good even without incorporating any other sophisticated TL techniques.
@@
Fig. 1 illustrates the flowchart of EA, using only one source domain. Most materials in this subsection are adapted from [13].
Figure 1: Illustration of EA and LA.
@@
Assume a domain (either source or target; different domains are processed identically) has $N$ EEG trials $\{X_{n}\}_{n=1}^{N}$, where $X_{n}\in\mathbb{R}^{c\times t}$, in which $c$ is the number of EEG channels, and $t$ the number of time-domain sampling points. EA first computes the symmetric and semi positive-definite reference matrix $\bar{R}\in\mathbb{R}^{c\times c}$ as
@@
2.2 Where to Place EA in the BCI TL Pipeline
As proposed in our previous research [14], EA should be placed between the temporal filtering block and the spatial filtering block. This subsection performs experiments to demonstrate that this placement indeed leads to better performance.
Figure 2: The optimal location to place EA in the BCI TL pipeline.
We used the motor imagery Dataset 1 from BCI Competition IV [^6] http://www.bbci.de/competition/iv/desc_1.html. in our experiment, which was also extensively used in our previous research [13, 17, 14]. It was recorded from seven healthy subjects using 59 EEG channels at 1000 Hz, and later downsampled to 100 Hz. Each subject performed two classes of motor imagery tasks (100 trials per class) in the calibration session, selected from three options: left hand, right hand, and both feet. Only data from Subjects 2, 3, 4, 5 and 7 were used, as their imagined classes were consistent (left hand and right hand). As in [13], each trial consisted of EEG data between $[0.5,3.5]$ seconds after the cue appearance, i.e., each trial was a matrix with dimensionality $59\times 300$.
We performed leave-one-subject-out cross-validation in TL: each time, we picked one subject as the test subject (target domain), and combined all trials from the remaining four subjects as the auxiliary data (source domain). The goal was to facilitate the classifier training for the target subject using the source data. The number of labeled calibration trials in the target domain, $M$, increased from 0 to 20, with step size 4. The $M$ trials were randomly selected in a continuous block, to avoid the block-design pitfall in BCIs [18]. The performance of different approaches was evaluated using the classification accuracy on the remaining $100-M$ target domain samples. The experiments were repeated 30 times to cope with randomness.
@@ 2.2 Where to Place EA in the BCI TL Pipeline
    No EA, as shown in Fig. 3.
    Figure 3: Different configurations of the TL pipeline for motor imagery based BCIs. (a) No EA; (b) placement of EA after both temporal filtering and spatial filtering.
@@
    The above TL pipeline did not re-reference the EEG channels. We also investigated if performing average re-referencing (AR) after temporal filtering can further improve the TL performance. This approach is denoted as TF-AR-RCSP-wAR.
@@
    The recommended placement, between temporal filtering and spatial filtering, as shown in Fig. 2.
@@
    The above TL pipeline did not re-reference the EEG channels. We also evaluated another slightly modified pipeline that performed AR immediately after temporal filtering, which is denoted as TF-AR-EA-RCSP-wAR.
@@
    Placing EA after both temporal filtering and spatial filtering, as shown in Fig. 3.
@@
    Again, the above TL pipeline did not re-reference the EEG channels. We also evaluated another slightly modified pipeline that performed AR immediately after temporal filtering, which is denoted as TF-AR-RCSP-EA-wAR.
Fig. 4 shows the TL performance of the six pipelines, averaged over 30 repeated runs. The last subfigure shows the TL performance, further averaged across the five subjects. We can observe that:
Figure 4: The motor imagery classification performance of six different TL pipelines.
@@
    Performing EA, not matter whether it was between temporal filtering and spatial filtering, or after both temporal filtering and spatial filtering, always improved the TL performance, i.e., TF-EA-RCSP-wAR and TF-RCSP-EA-wAR always outperformed TF-RCSP-wAR, and TF-AR-EA-RCSP-wAR and TF-AR-RCSP-EA-wAR always outperformed TF-AR-RCSP-wAR. This is because EA always reduces the data distribution discrepancies among the subjects, no matte where it is placed, and reducing these discrepancies always benefit TL.
@@
    Performing EA between temporal filtering and spatial filtering almost always achieved better performance than performing EA after both temporal filtering and spatial filtering, i.e., on average TF-EA-RCSP-wAR outperformed TF-RCSP-EA-wAR, and TF-AR-EA-RCSP-wAR outperformed TF-AR-RCSP-EA-wAR. This suggests that EA should enter the TL pipeline early to achieve maximum benefits.
@@
    Performing average re-referencing often led to performance improvement, i.e., TF-AR-RCSP-wAR outperformed TF-RCSP-wAR for four of the five subjects and on average, TF-AR-EA-RCSP-wAR slightly outperformed TF-EA-RCSP-wAR on average, and TF-AR-RCSP-EA-wAR always outperformed TF-RCSP-EA-wAR. This is because average referencing helps reduce the magnitude differences between EEGs from different subjects.
@@
In summary, we confirmed through experiments that EA should be placed between temporal filtering and spatial filtering in motor imagery based BCIs, as in Fig. 2, for maximum benefits; and, performing average referencing immediately after temporal filtering could further improve the performance.
2.3 Applications of EA

### v2

@@ Abstract
Due to large intra-subject and inter-subject variabilities of electroencephalogram (EEG) signals, EEG-based brain-computer interfaces (BCIs) usually need subject-specific calibration to tailor the decoding algorithm for each new subject, which is time-consuming and user-unfriendly, hindering their real-world applications. Transfer learning (TL) has been extensively used to expedite the calibration, by making use of EEG data from other subjects/sessions. An important consideration in TL for EEG-based BCIs is to reduce the data distribution discrepancies among different subjects/sessions, to avoid negative transfer. Euclidean alignment (EA) was proposed in 2020 to address this challenge. Numerous experiments from 13 different BCI paradigms demonstrated its effectiveness and efficiency. This paper revisits EA, explaining its procedure and correct usage, introducing its applications and extensions, and pointing out potential new research directions. It should be very helpful to BCI researchers, especially those who are working on EEG signal decoding.
Index Terms:
Brain-computer interface, EEG, Euclidean alignment, label alignment, transfer learning
I Introduction
@@
According to the input signals, BCIs can be broadly classified into invasive ones and non-invasive ones. The former use surgery to implant electrodes into the brain, which are mainly used for patients. Non-invasive BCIs do not need surgery, and hence are more convenient and preferred by able-bodied users. Electroencephalogram (EEG) is the most frequently used input signal for non-invasive BCIs.
@@
EEG signals exhibit both intra-subject variability (non-stationarity), e.g., the same subject may have different EEG responses to the same stimulus in different sessions, and inter-subject variability (individual differences), i.e., different subjects may have significantly different EEG responses to the same stimulus. As a result, it is very challenging, if not impossible, to design an EEG-based non-invasive BCI system that is plug-and-play and fits all subjects. Usually, for each new subject, we need to collect some subject-specific calibration data to tailor the parameters of the decoding algorithm, which is time-consuming and user-unfriendly.
@@
Transfer learning (TL), which utilizes data/knowledge from other subjects (sessions) to facilitate the calibration for a new subject (session), has been extensively used in EEG-based BCIs to expedite the calibration [11]. An important consideration in TL is how to reduce the EEG data distribution discrepancies among different subjects (sessions) to avoid negative transfer [12].
@@
Euclidean Alignment (EA) [13] was proposed in 2020 to address this challenge. It uses two simple, efficient and closed-form formulas to align the raw EEG trials from different subjects/sessions, greatly facilitating TL [14]. It also inspires some other EEG data alignment approaches, e.g., label alignment (LA) [15]. The EA paper has been cited 283 times in WoS and 405 times in Google Scholar (as of May 3, 2025), ranking 2/344 of all papers published in IEEE Transactions on Biomedical Engineering [^2] https://webofscience.clarivate.cn/wos/alldb/summary/fd21b6a6-bdf1-4dcd-bae6-45d4041b195f-014738bfd9/times-cited-descending/1 (TMBE) in 2020, and 3/1872 of all papers published in IEEE TBME [^3] https://webofscience.clarivate.cn/wos/alldb/summary/5e219ca8-1532-4637-a27d-6c70f9cbbca2-014738a148/times-cited-descending/1 between 2020 and 2025. It was one of the only two Highly Cited Papers of IEEE TBME in 2020. It was also the number one contributor to the 2022 impact factor (4.6) of IEEE TBME [^4] https://jcr.clarivate.com/jcr-jp/journal-profile?journal=IEEE%20T%20BIO-MED%20ENG&year=2022&fromPage=%2Fjcr%2Fhome, among all 695 citable items published in 2020 and 2021.
@@
After five years of the publication of the EA paper, we now have deeper and more comprehensive understanding on how to properly use EA. Its applicable BCI paradigms have also been greatly expanded. This paper revisits the EA approach, explaining its procedure and correct usage, introducing its applications and extension, and pointing out potential new research directions. It should be very helpful to BCI researchers, especially those who are working on EEG signal decoding.
@@
The remainder of this paper is organized as follows: Section II introduces the procedure, correct usage, and applications of EA. Section III introduces an extension of EA, and some related EEG data alignment approaches. Section IV points out potential future research directions. Finally, Section V draws conclusions.
II Euclidean Alignment (EA)
@@
This section introduces the EA approach [13]. It was inspired by Riemannian alignment [16] proposed in 2018 in IEEE TBME, which has been cited 272 times in WoS [^5] https://webofscience.clarivate.cn/wos/alldb/summary/bbb3db2d-ab56-4291-8b2e-884a88511745-015ed46b9e/times-cited-descending/1 and 386 times in Google Scholar (as of may 3, 2025) and is one of the five Highly Cited Papers of IEEE TBME in 2018. Riemannian alignment has demonstrated promising performance in seven different BCI paradigms, as shown in Table I.
TABLE I: Classification accuracy improved by Riemannian alignment in 7 different BCI paradigms. Note that different publications generally used different datasets.
No. BCI Paradigm Classifier Accuracy (%) Reference
w/o RA w/ RA Improvement
1 Motor Imagery MDM 38.84 63.76 24.92 Zanini et al., IEEE TBME 2018 [16]
GM-4 41.58 64.18 22.60
MDM 50.14 66.21 16.07 He and Wu, IEEE SMC 2019[17]
MDM 64.81 66.82 2.01 Kumar et al., BCI 2019 [18]
FgMDM 68.94 71.10 2.16
MDM 57.18 70.91 13.73 He and Wu, IEEE TBME 2020 [13]
FBCSP-LDA 69.00 70.00 1.00 Xu et al., FHN 2020 [19]
ShallowNet 71.00 72.00 1.00
EEGNet 70.00 72.00 2.00
EEGNet 69.70 71.40 1.70 Xu et al., JNE 2021 [20]
EEGNetv4 67.40 70.30 2.90
MDM 57.02 72.45 15.43 Zhang et al., CMPB 2021 [21]
MDM 57.18 70.91 13.73 Liang et al., IEEE IROS 2021 [22]
MDM 56.00 70.00 14.00 Li et al., IEEE Access 2021 [23]
MDM 57.18 70.91 13.73 Liang et al., ICIEA 2022 [24]
EEGNet 53.45 57.93 4.48 Zoumpourlis and Patras, BCI 2022 [25]
MDM 32.75 50.93 18.18 Xia et al., IEEE TBME 2022 [26]
MDM 75.70 76.73 1.03 Pan et al., JNE 2023 [27]
MDM 57.18 70.91 13.73 Liang et al., BSPC 2024 [28]
BaseNet 57.38 63.64 6.26 Wimpff et al., BCI 2024 [29]
EEGNet 72.18 76.42 4.24 Junqueira et al., JNE 2024 [30]
DeepNet 69.27 76.92 7.65
ShallowNet 69.99 74.23 4.24
MDM 57.18 70.91 13.73 Chu et al., JNE 2024 [31]
MDM 54.36 73.57 19.21 Jiang et al., ICASSP 2024 [32]
LDA 52.07 70.64 18.57
CSP-MDM 61.78 87.93 26.15
CSP-LDA 65.71 83.64 17.93
JF-MDM 82.36 94.3 11.94
JF-LDA 80.21 94.36 14.15
2 Event Related Potential MDM 51.43 87.48 36.05 Zanini et al., IEEE TBME 2018[16]
GM-4 45.99 86.18 40.19
3 Mental Imagery MDM 58.33 69.71 11.38 Kumar et al., BCI 2019 [18]
FgMDM 58.50 70.35 11.85
4
Steady-State
Visual Evoked
Potential
MDM 69.30 73.00 3.70 Rodrigues et al., IEEE TBME 2019 [33]
5
Motor+Speech
Imagery
MDM 51.32 54.82 3.50 Zhan et al., CBM 2022 [34]
6
Mindfulness
Meditation
EEGNet 54.88 69.93 15.05 Karaiskou et al., MetroXRAINE 2023[35]
7
Natural Hand
Movement
Decoding
MDM 53.33 62.99 9.66 Xue et al., Measurement 2025 [36]
II-A The EA Algorithm
@@
The goal of EA is to make the marginal distributions of EEG data from different subjects/sessions more similar, i.e., to reduce the inter-subject/session EEG data discrepancies, facilitating TL. In fact, after EA, often the EEG data from auxiliary subjects/sessions (called source domains in TL) can be combined directly with the limited amount of calibration data from the new subject/session (called target domain in TL) to train a classifier; the performance is usually quite good even without incorporating any other sophisticated TL techniques.
@@
Figure 1 illustrates the flowchart of EA, using only one source domain. Most materials in this subsection are adapted from [13].
Figure 1: Illustration of Euclidean alignment and label alignment. Euclidean alignment aligns the marginal probability distributions of the source and target domains, whereas label alignment aligns the class conditional probability distributions of different domains.
@@
Assume a domain (either source or target; different domains are processed individually and identically) has $N$ EEG trials $\{X_{n}\}_{n=1}^{N}$, where $X_{n}\in\mathbb{R}^{c\times t}$, in which $c$ is the number of EEG channels, and $t$ the number of time-domain sampling points. EA first computes the symmetric and semi positive-definite reference matrix $\bar{R}\in\mathbb{R}^{c\times c}$ as
@@
II-B Where to Place EA in the Traditional BCI TL Pipeline
As proposed in our previous research [14] and illustrated in Figure 2, when traditional machine learning classifiers with manually extracted features are used, EA should be placed between the temporal filtering block and the spatial filtering block. This subsection performs experiments to demonstrate that this placement indeed leads to better performance.
Figure 2: The optimal location to perform Euclidean alignment in the traditional BCI transfer learning pipeline (between temporal filtering and spatial filtering).
We used the motor imagery Dataset 1 from BCI Competition IV [^6] http://www.bbci.de/competition/iv/desc_1.html. in our experiment, which was also extensively used in our previous research [13, 15, 14]. It was recorded from seven healthy subjects using 59 EEG channels at 1000 Hz, and later downsampled to 100 Hz. Each subject performed two classes of motor imagery tasks (100 trials per class) in the calibration session, selected from three options: left hand, right hand, and both feet. Only data from Subjects 2, 3, 4, 5 and 7 were used, as their imagined classes were consistent (they all imagined the movements of left hand and right hand, whereas Subjects 1 and 6 imagined the movements of left hand and both feet). As in [13], each trial consisted of EEG data between $[0.5,3.5]$ seconds after the cue appearance, i.e., each trial was a matrix with dimensionality $59\times 300$.
We performed leave-one-subject-out cross-validation in TL: each time, we picked one subject as the test subject (target domain), and combined all trials from the remaining four subjects as the auxiliary data (source domain). The goal was to facilitate the classifier training for the target subject using the source data. The number of labeled calibration trials in the target domain, $M$, increased from 0 to 20, with a step size of 4. The $M$ trials were randomly selected in a continuous block, to avoid the block-design pitfall in BCIs [37]. The performance of different approaches was evaluated using the classification accuracy on the remaining $100-M$ target domain samples.
@@ II-B Where to Place EA in the Traditional BCI TL Pipeline
    No EA, as shown in Figure 3.
    Figure 3: Different configurations of the transfer learning pipeline for motor imagery based BCIs. (a) No Euclidian alignment; (b) performing Euclidean alignment after both temporal filtering and spatial filtering.
@@
    The above TL pipeline does not re-reference the EEG channels. We also investigated if performing common average referencing (CAR) after temporal filtering can further improve the TL performance. This approach is denoted as TF-CAR-RCSP-wAR.
@@
    The recommended placement, between temporal filtering and spatial filtering, as shown in Figure 2.
@@
    The above TL pipeline did not re-reference the EEG channels. We also evaluated another slightly modified pipeline that performed CAR immediately after temporal filtering, which is denoted as TF-CAR-EA-RCSP-wAR.
@@
    Placing EA after both temporal filtering and spatial filtering, as shown in Figure 3.
@@
    Again, the above TL pipeline did not re-reference the EEG channels. We also evaluated another slightly modified pipeline that performed CAR immediately after temporal filtering, which is denoted as TF-CAR-RCSP-EA-wAR.
Figure 4 shows the TL performance of the six pipelines, averaged over 30 repeated runs to accommodate randomness. The last subfigure shows the overall TL performance, further averaged across the five subjects. We can observe that:
Figure 4: The motor imagery classification performance of six different transfer learning pipelines. TF: Temporal filtering; RCSP: Regularized common spatial pattern; wAR: weighted adaptation regularization; CAR: Common average referencing; EA: Euclidean alignment.
@@
    Performing EA, no matter whether it was between temporal filtering and spatial filtering, or after both temporal filtering and spatial filtering, always improved the TL performance, i.e., TF-EA-RCSP-wAR and TF-RCSP-EA-wAR always outperformed TF-RCSP-wAR, and TF-CAR-EA-RCSP-wAR and TF-CAR-RCSP-EA-wAR always outperformed TF-CAR-RCSP-wAR. This is because EA always reduces the marginal EEG data distribution discrepancies among the subjects, no matter where it is placed, and reducing these discrepancies always benefits TL.
@@
    Performing EA between temporal filtering and spatial filtering almost always achieved better performance than performing EA after both temporal filtering and spatial filtering, i.e., on average TF-EA-RCSP-wAR outperformed TF-RCSP-EA-wAR, and TF-CAR-EA-RCSP-wAR outperformed TF-CAR-RCSP-EA-wAR.
    EA should be placed after temporal filtering, because temporal filtering improves the EEG signal quality by smoothing it; particularly, it suppresses large EEG signal values, which may be outliers. This in turn improves the quality of the covariance matrices, and subsequently the EA reference matrix $\bar{R}$ and transformation matrix $\bar{R}^{-1/2}$.
    EA should be placed before spatial filtering, because RCSP utilizes EEG trials from both source and target subjects; EA reduces the marginal distribution discrepancies of EEG trials from different subjects, so that RCSP can identify spatial filters that work well for all subjects.
@@
    Performing CAR often led to performance improvement, i.e., TF-CAR-RCSP-wAR significantly outperformed TF-RCSP-wAR for four of the five subjects and on average, TF-CAR-EA-RCSP-wAR slightly outperformed TF-EA-RCSP-wAR on average, and TF-CAR-RCSP-EA-wAR always significantly outperformed TF-RCSP-EA-wAR. This is because CAR helps reduce the magnitude differences between EEG trials from different subjects.
  4. 4.
    For the best performing TF-CAR-EA-RCSP-wAR, the average binary classification accuracy between left hand and right hand for the five subjects, without any labeled target-specific calibration data (i.e., $M=0$), was 81.60%, and the classification accuracy increased very slowly with $M$ (e.g., the accuracy was 83.96% when $M=20$). These suggest that EA exploits rich information from the source subjects, and enables plug-and-play motor imagery classification for a new subject.
@@
In summary, we confirmed through experiments that EA should be placed between temporal filtering and spatial filtering in traditional TL pipeline for motor imagery based BCIs, as in Figure 2; and, performing CAR immediately after temporal filtering could further improve the performance.
II-C Where to Place EA in the BCI Deep TL Pipeline
When end-to-end deep learning approaches, e.g., EEGNet [40], are used, EA should also be placed between the temporal filtering block and the deep learning module, as illustrated in Figure 5. This subsection performs experiments to demonstrate that this placement indeed leads to better performance.
Figure 5: The optimal location to perform Euclidean alignment in the BCI deep transfer learning pipeline (after temporal filtering).
The dataset in the previous subsection was used again. We chose the popular EEGNet [40] as the deep learning model, and compared three different configurations:
  1. 1.
    EEGNet-only, in which no band-pass temporal filtering and no EA were performed, i.e., the raw EEG signals were epoched into 3-second trials, and trials from all subjects were combined and fed directly into EEGNet.
  2. 2.
    EA-EEGNet, in which EA was performed without temporal filtering, i.e., the raw EEG signals from each subject were epoched into 3-second trials and aligned by EA, and then trials from all subjects were combined and fed directly into EEGNet.
  3. 3.
    TF-EA-EEGNet, as shown in Figure 5, in which the raw EEG signals from each subject were [8,30] Hz band-pass filtered, epoched into 3-second trials, aligned by EA, and then trials from all subjects were combined and fed into EEGNet.
Due to much higher computational cost of EEGNet than traditional machine learning approaches, we performed only plug-and-play cross-subject validations, i.e., the target subject had no labeled calibration trials at all ($M=0$). The experiments for each subject were repeated three times to accommodate randomness, and the average results are reported in Table II. Both EEGNet-only and EA-EEGNet had near-random binary classification accuracies, but TF-EA-EEGNet performed much better, suggesting the necessity to apply temporal filtering explicitly before EA in deep TL.
TABLE II: Classification accuracies (%) of the three deep transfer learning configurations.
Subject 1 2 3 4 5 Average
EEGNet-only 49.83 50.00 48.00 49.00 49.83 49.33
EA-EEGNet 52.67 46.83 43.17 46.83 50.50 48.00
TF-EA-EEGNet 70.17 59.83 59.50 91.50 66.33 66.33
II-D Consistency of the EA Transformation Matrix Among Different Subjects
The original EEG trials from different subjects have consistent channel configurations before alignment, because they use the same EEG headset. For example, for the MI trials in our experiments, the first channel was always AF3, the second channel was AF4, and the 59th channel was O2. It’s important to maintain this channel consistency after EA, as channel location shifting may impact the subsequent classification performance.
Figure 6 uses heatmap to visualize the elements of the EA transformation matrix $\bar{R}^{-1/2}\in\mathbb{R}^{59\times 59}$ for the five subjects. There is a clear diagonal pattern, indicating that the original $i$-th channel in $X_{n}$ contributes the most to the transformed $i$-th channel in $\tilde{X}_{n}$. Hence, the channel consistency among different subjects is maintained after EA, although their transformation matrices are computed individually. This is another reason why EA achieves promising performance.
Figure 6: Visualization of the EA transformation matrix $\bar{R}^{-1/2}$.
II-E Applications of EA